OpenStudy (anonymous):

If x cannot equal 3 or -4, and 7x-21/x-3 = x^2-x-20/x+4, what is the value of x? Choices are: -12 4 7 10 12

OpenStudy (anonymous):

\[\large \frac{7(x-3)}{(x-3)} = \frac{(x-5)(x+4)}{(x+4)}\] Cancel what you can cancel here..

OpenStudy (anonymous):

(x-3)(X^2-x-20) = (7x-21)(x+4) (x-3)(x-5)(x+4) = 7(x-3)(x+4) Cancel(x+4) and (x-3) from both sides you will get: x-5 = 7 So x = 12

OpenStudy (anonymous):

Thank you!!

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

i looked at my answer key but it says it should be -12..